OpenStudy (anonymous):

What is the slope of the line with points (7, 6) and (–2, 2)?

OpenStudy (anonymous):

m=4/9 if i remember hw to do this correctly

OpenStudy (anonymous):

kly_aqui.no is right.

OpenStudy (anonymous):

\[(x _{1}-x _{2})/(y _{1}-y _{2})\] that would be your equation

OpenStudy (anonymous):

yep

OpenStudy (skullpatrol):

@kly_aqui.no You have the formula UP SIDE down

OpenStudy (anonymous):

woops, sorry. you're right. im not exactly at my best concentration right now

OpenStudy (skullpatrol):

$$\huge m = \frac{y_2-y_1}{x_2-x_1}$$

OpenStudy (anonymous):

much better, thanks

OpenStudy (skullpatrol):

$$\huge \text{slope =}{ \frac{rise}{run}}$$